OpenStudy (anonymous):

(x + y)(x - y) i got as an answer x^2-xy?

OpenStudy (anonymous):

you almost had it right just change -xy to -y^2

OpenStudy (anonymous):

(x+y)(x-y) = x^2-y^2

OpenStudy (anonymous):

x^2-xy+yx-y^2 x^2-y^2

OpenStudy (anonymous):

Another helpful rule...\[(a+b)(a-b)=a^{2}-b^{2}\] This is called the difference of squares formula